# Creating a Historical Migration Dataset from Finnish Church Records, 1800–1920

### v1

@@ 1 Introduction
The contribution of this work is thus two-fold. Firstly, we develop and evaluate a pipeline for the extraction of data from the Finnish church migration records. Our pipeline applies machine learning throughout the workflow — image pre-processing, table structure detection, and text recognition — to convert data into a format suitable for research. Although most likely not constituting a universal solution, the individual steps and their evaluation will add to the spectrum of techniques available for historical tabular data extraction. Further, with machine learning playing an increasing role in cultural heritage research (Fiorucci et al., 2020), the methods developed here add to our understanding of the possibilities and limitations of these techniques to support research across the humanities and social sciences by linking archival sources with computational approaches.
Our second contribution is a dataset obtained using the pipeline we develop, capturing detailed information on individuals and families who moved within Finland between 1800 and 1920. In total, the dataset constitutes over six million entries extracted from approximately 200,000 images of church migration records. These records provide a basis for studying internal migration trends and examining the social, economic, and health-related effects of mobility in historical contexts. Furthermore, once the dataset has been linked with other historical sources — such as records of births, deaths, and diseases — it can contribute to research on how mobility contributed to the spread of infectious diseases like smallpox, measles, and pertussis (Briga et al., 2022; Ketola et al., 2021; Nitsch et al., 2025). It also supports research on broader topics such as economic development, urbanization, and transformations in rural communities. If combined with census data or tax registers, the dataset can enable longitudinal studies of demographic change. The dataset is openly available, aiming to support future research into population movement, demographic change, and regional development using historical records. It can also support reconstructing individual life courses and family histories.
@@ 2 Data
The number of columns and the content present in the migration tables vary; however, the most common information is presented in Table 2 and Figure 4 shows one complete entry of an out-migration record. Occasionally, the same field may be split into two or more columns, for example, the date may appear as one field or be divided into separate columns for the month and the day of the month. Typically, one person is recorded on a single row of the table, with other members of the same family listed on subsequent rows. In older books, all members of the same household are sometimes grouped together in a single record entry. The language of the records also varies between books, with some of the records written in Finnish and others in Swedish because both are used as administrative languages in Finland. Furthermore, many parishes have Finnish and Swedish names (e.g. Helsinki – Helsingfors), and the orthography of these names in historical data may differ or include abbreviations (e.g. Helsingfors – H:fors). These, and other inconsistencies in data recording practices, as well as missing or incomplete entries, further complicate the extraction process.
@@ 2.1 Manual annotation for model training and evaluation
Annotation type Count of Train Dev Test Total
De-skew key points images 900 190 200 1,290
Table structure images 1,252 188 192 1,632
Cell type images/cells 230/47K 47/14K 46/16K 323/77K
Text recognition images/cells 41/1,947 39/2,277 80/4,224
Year recognition images 1,026 188 192 1,326
Table 3: Summary of manually annotated data for the different stages of the pipeline, divided into training, development and test sections.
@@ 3.1 Text recognition pipeline for tabular data
Figure 5 outlines the main phases of the workflow, which relies on OCR and handwritten text recognition (HTR) techniques. During the preprocessing phase, the system converts image files into a uniform format and improves their quality for subsequent steps. Preprocessing includes tasks such as image de-skewing, resizing, normalization, noise reduction, and binarization.
@@
Text recognition involves two main steps: document layout analysis and recognition. Layout analysis divides the image into smaller sections (e.g., text regions, tables, and lines) for further processing. Character and word recognition occur during the text recognition phase. In this step, the system identifies characters, and an XML (extensible markup language) file or similar format is generated, which includes references to object locations in the original image file and the recognized text.
@@ 3.3 Table structure detection
YOLO is used for Table Detection for its speed and accuracy (Deng et al., 2024; Huang et al., 2019; Ultralytics, 2024). It processes an entire image in a single forward pass, making it efficient for detecting multiple tables in a document. The model is trained to detect entire tables as well as individual table cells (Figure 8). Some cells may be missed during prediction. To address this, density-based clustering (DBSCAN) is applied as a postprocessing step. The detected cell borders serve as elements to be clustered—left and right borders for columns, and top and bottom borders for rows. This helps infer missing cells and reconstruct the complete table structure efficiently.
@@ 3.5 Text recognition
For this task, we use a handwriting recognition model (HTR) model trained by the National Archives of Finland [^1] https://huggingface.co/Kansallisarkisto/tablecell-htr. The model is based on the TrOCR architecture introduced by Li et al. (2023), a transformer-based encoder-decoder model consisting of an image encoder and a text decoder. While the original TrOCR model was trained on English data, the National Archives of Finland fine-tuned it for handwritten, historical Finnish and Swedish. The fine-tuning dataset included text lines from a variety of sources spanning the 17th to 20th centuries, as well as tabular data to improve model’s performance on table cells.
@@ 4 Results and error analysis
Among these, errors (1) and (3) can potentially be addressed through post-processing, while error (2) would require improvements to the table recognition component.
@@ 5 Case study: Migrations of Elimäki
All six books use preprinted layouts, which allowed us to rely on our metadata annotations to determine whether a page contains in- or out-migration records (in books with mixed records, the left page corresponds to in-migration and the right to out-migration), and to identify the expected column for the parish name. [^2] We note that one reason we selected Elimäki as our case study parish is that all its books use preprinted forms, which simplifies the extraction process. Section 6 discusses our plans for extending this approach to books with handdrawn tables.
@@
This process resulted in over 3,000 unique parish names. However, this number includes both minor orthographic variations in the original writings as well as cases where the text recognition model misidentified some of the letters. To remove noise and normalize the names, we first applied LLM-based cleaning, where we prompted OpenAI’s GPT-4o-mini333gpt-4o-mini-2024-07-18 to map the predicted parish names to a given list of known Finnish parishes. Since the relationship between the predicted name and the actual parish was not always straightforward, the model was instructed to flag uncertain cases. While this approach resolved many of the simple cases, we manually reviewed the LLM output and further standardized it by matching each name to the known list of parishes. After standardization, 66% of the 3,363 parish name spellings could be linked to a specific parish.
@@ 5.1 Case study results
This dataset enables us to quantify the departures from and arrivals to Elimäki across both time and space. For instance, it is possible to quantify annual departures (Mean = 135 departures ±$\pm 59$ s.d., Figure 10) and arrivals to the parish (Mean = 100 ±$\pm 41$ s.d., Figure 10). It is worth noting that there are no records of arrivals to Elimäki for the years 1914 and 1915. This is due to the year identification failing in few pages, and the records from these years are merged with records from 1916 and 1917.
@@
Spatially, the initial overview of the destinations and the origins of migration shows that most movements were local (see Figure 11 and Supplementary videos available on the project repository [^4] https://github.com/TurkuNLP/finnish-migration-data for a yearly overview).
@@ 6 Conclusions and future work
This work demonstrates that large-scale, automated extraction of structured data from handwritten historical records is feasible, even though it remains technically challenging. Using digitized Finnish moving records, we created a dataset containing over six million entries from approximately 200,000 images dated between the 19th and early 20th centuries. These entries document individual and family migration between parishes and provide a valuable source for demographic and historical analysis. The resulting dataset has a potential to support studies of population movement, urbanization, family structures, and disease transmission in preindustrial Finland. In a case study based on the Elimäki parish, we illustrated how migration volumes and patterns can be reconstructed locally.
@@
Our pipeline combines deep learning methods for preprocessing, layout analysis, table detection, cell classification, and handwriting recognition. Its modular structure allows each stage to be improved independently. While we have already demonstrated the extraction output to be sufficient for certain quantitative applications, there are still challenges that need to be addressed in future development. First, we aim to enhance the extraction of document metadata, such as table type, table titles, and column headers. Additional metadata is needed for conducting studies similar to our case study, especially when working with handdrawn books, where the expected columns on a page are not known in advance. To address this, we intend to use LLM-based column header inference, where a language model will inspect the extracted data table and infer common column headers (e.g. date, name, and parish).
@@
Second, we are developing LLM-based postprocessing tools to standardize variations in the extracted data. This will help to normalize e.g. names, places, and dates including both orthographic variation as well as misread letters into standardized spelling, as well as infer missing values where contextual patterns in the data suggest repetition or shared information.

### v2

@@ 1 Introduction
The contribution of this work is thus two-fold. Firstly, we develop and evaluate a pipeline for the extraction of data from the Finnish church migration records. Our pipeline applies machine learning throughout the workflow — image pre-processing, table structure detection, and text recognition — to convert data into a format suitable for research. Although most likely not constituting a universal solution, the individual steps and their evaluation will add to the spectrum of techniques available for historical tabular data extraction. Further, with machine learning playing an increasing role in cultural heritage research (Fiorucci et al., 2020), the methods developed here add to our understanding of the possibilities and limitations of these techniques to support research across the humanities and social sciences by linking archival sources with computational approaches. We also identify key limitations of the current pipeline, and in the future work section discuss our ideas for post-processing to address them.
Our second contribution is a dataset obtained using the pipeline we developed, capturing detailed information on individuals and families who moved within Finland between 1800 and 1920. This dataset is openly available through Zenodo [^1] https://zenodo.org/records/15606656, ensuring transparency, reproducibility, and accessibility for future research. In total, the dataset constitutes over six million entries extracted from approximately 200,000 images of church migration records. These records provide a basis for studying internal migration trends and examining the social, economic, and health-related effects of mobility in historical contexts. Furthermore, once the dataset has been linked with other historical sources — such as records of births, deaths, and diseases — it can contribute to research on how mobility contributed to the spread of infectious diseases like smallpox, measles, and pertussis (Briga et al., 2022; Ketola et al., 2021; Nitsch et al., 2025). It also supports research on broader topics such as economic development, urbanization, and transformations in rural communities. If combined with census data or tax registers, the dataset can enable longitudinal studies of demographic change. The dataset is openly available, aiming to support future research into population movement, demographic change, and regional development using historical records. It can also support reconstructing individual life courses and family histories.
@@ 2 Data
The number of columns and the content present in the migration tables vary; however, the most common information is presented in Table 2 and Figure 4 shows one complete entry of an out-migration record. Occasionally, the same field may be split into two or more columns, for example, the date may appear as one field or be divided into separate columns for the month and the day of the month. Typically, one person is recorded on a single row of the table, with other members of the same family listed on subsequent rows. In older books, all members of the same household are sometimes grouped together in a single record entry. The language of the records varies across books. Some entries are written in Finnish, others in Swedish, as both have long served as administrative languages in Finland. Many parishes have names in both languages, such as Helsinki and Helsingfors. In historical sources, these names may appear with altered spelling or abbreviations, for example Helsingfors written as H:fors.
Many of the place names remain uncertain. One reason is that Swedish-speaking clerks often wrote Finnish names phonetically. This led to unusual or hybrid forms. Spelling was not yet standardized in the 19th century, so a single place could appear under several names. For instance, Mäntyharju might be recorded as Mändyharju, Menduharju, or Menduhariu. Local usage adds more variation. People often used names that differ from maps or official lists. A location could have one name as a farm, another as a village, and yet another as a district. These naming practices are well documented in Finnish place name research (Paikkala et al., 2007).
Correctly resolving place names requires not only accurate recognition but also historical understanding of how those names were written and used over time. The task of recognizing and grounding historical place names is a significant area of research. For an overview of current methods and benchmarks, see the recent review by Ehrmann et al. (2023), and for an example of a complete pipeline, refer to Pontes et al. (2022). Additionally, the well-known Pelagios Network (Barker et al., 2021) supports community-driven linked open data initiatives for historical places, providing both annotation tools and infrastructure for name grounding.
@@ 2.1 Manual annotation for model training and evaluation
Annotation type Train Dev Test Total
Images Cells Images Cells Images Cells Images Cells
De-skew key points 900 190 200 1,290
Table structure 1,252 188 192 1,632
Cell type 230 47,000 47 14,000 46 16,000 323 77,000
Text recognition 41 1,947 39 2,277 80 4,224
Year recognition 1,026 188 192 1,326
Table 3: Summary of manually annotated data for different stages of the pipeline, divided into training, development, and test sets. Image and cell counts are shown separately.
@@ 3.1 Text recognition pipeline for tabular data
Figure 5 outlines the main phases of the workflow, which relies on OCR and handwritten text recognition (HTR) techniques. During the preprocessing phase, the system converts image files into a uniform format and improves their quality for subsequent steps. Preprocessing includes tasks such as image de-skewing, resizing, and techniques like normalization (adjusting pixel intensity values to a standard range), noise reduction (removing unwanted visual artifacts), and binarization (converting the image to black and white to separate text from background).
@@
Text recognition involves two main steps: document layout analysis and recognition. Layout analysis divides the image into smaller sections (e.g., text regions, tables, and lines) for further processing. Character and word recognition occur during the text recognition phase. In this step, the system identifies individual characters, and the output is stored in a structured format—specifically, the PAGE XML 2013 schema. This format, supported by Transkribus, our primary annotation tool, includes references to object locations in the original image file along with the corresponding recognized text.
@@ 3.3 Table structure detection
YOLO is used for Table Detection for its speed and accuracy. It processes an entire image in a single forward pass, making it efficient for detecting multiple tables in a document. The model is trained to detect entire tables as well as individual table cells (Figure 8). Some cells may be missed during prediction. To address this, density-based clustering (DBSCAN) is applied as a postprocessing step. The detected cell borders serve as elements to be clustered—left and right borders for columns, and top and bottom borders for rows. This helps infer missing cells and reconstruct the complete table structure efficiently.
@@ 3.5 Text recognition
For this task, we use a handwriting recognition (HTR) model trained by the National Archives of Finland (Kansallisarkisto, 2024). The model is based on the TrOCR architecture introduced by Li et al. (2023), a transformer-based encoder-decoder model consisting of an image encoder and a text decoder. While the original TrOCR model was trained on English data, the National Archives of Finland fine-tuned it for handwritten, historical Finnish and Swedish. The fine-tuning dataset included more than 700,000 transcribed text lines from a wide range of historical sources spanning the 17th to 20th centuries. In addition to narrative texts, the training data also incorporated tabular materials to enhance the model’s performance on structured documents such as tables and registers.
As the text recognition model is readily available, we have directly integrated it into our processing pipeline without further modification. Since our primary focus is on developing methods for structuring and analyzing the extracted data, the handwriting recognition component itself is not the central topic of this article.
@@ 4 Results and error analysis
A visual inspection of the table structure detection results confirms that the method performs consistently across both printed and hand-drawn Finnish church records. Nevertheless, several recurring error types were identified. The most frequent issue involves two-page tables, which are often detected as two separate tables, rather than a single continuous one. On sparsely populated pages, overlapping detections may occur when both a partial and full-page table are recognized. Row detection errors include the false detection of empty rows, missed final rows, and misplaced boundaries when a cell contains multiple lines of text. Column detection errors include missed edge columns, incorrect splitting of wide columns, and merging of adjacent narrow columns. These issues are often linked to irregular layouts, weak or faded lines, and variable content density—factors commonly found in historical documents.
To complement the quantitative evaluation, we have created a supplementary document illustrating these typical errors through annotated examples. This material is available via the project’s GitHub repository and provides further insight into the method’s current limitations and potential areas for improvement.
@@
Among these, errors (1) and (3) can potentially be addressed through post-processing, while error (2) would require improvements to the table recognition component. In future work, we will investigate post-processing approaches to merge the double-page tables into one, as well as realign the columns to known headers in cases where the predicted table has too few or too many columns. The realignment will require understanding which columns are extra and which are missing, which should be relatively straightforward if the difference is small but a lot harder with larger discrepancies. In the future work section (Section 6), we discuss other ideas for post-processing.
@@ 5 Case study: Migrations of Elimäki
All six books use preprinted layouts, which allowed us to rely on our metadata annotations to determine whether a page contains in- or out-migration records (in books with mixed records, the left page corresponds to in-migration and the right to out-migration), and to identify the expected column for the parish name. We note that one reason we selected Elimäki as our case study parish is that all its books use preprinted forms, which simplifies the extraction process. In the case of handdrawn formats, information about migration direction exists for most of the books with only a fraction having unknown direction, but we are currently lacking the information of which column contains the migration parish information. To extend the case study to include also handdrawn books, we need to infer this information from the data. Section 6 discusses our plans for extending this approach to books with handdrawn tables.
@@
This process resulted in over 3,000 unique parish names. The relationship between the predicted name and the actual parish is not always obvious for the reasons outlined in Section 2 related to the lack of systematicity in the records with respect to orthography and language, as well as failures of the text recognition model to resolve the handwriting correctly. We compared the list of unique parish names, as extracted by the text recognition model for the Elimäki test parish, to the list of known parishes maintained by the Finland’s Family History Association. We sanitized the raw records by removing leading and trailing non-alphabetic characters, as well as split the Finnish and Swedish parish names as their separate records. Table 11 summarizes the number of extracted parish names that have at least one potential matching record at different edit distances (edit distance is the number of character insertions, substitions, or deletions needed to match the strings). Only 8% can be paired directly with a known parish name, and only 23% has at least one candidate match at edit distance 0 or 1. Matching the parish names is therefore not a straightforward process.
$d=0$ $d\leq 1$ $d\leq 2$ $d\leq 3$ $d\leq 4$
8% 23% 41% 60% 72%
Table 11: Proportion of extracted parish names in the Elimäki books for which a known parish name can be found at edit distance of at most $d$.
To remove noise and normalize the names, we first applied LLM-based cleaning, where we prompted OpenAI’s GPT-4o-mini222gpt-4o-mini-2024-07-18 to map the predicted parish names to a given list of known Finnish parishes. While this approach resolved many of the simple cases, we manually reviewed the LLM output and further standardized it by matching each name to the known list of parishes. After standardization, 66% of the 3,363 parish name spellings could be linked to a specific parish.
@@ 5.1 Case study results
This dataset enables us to quantify the departures from and arrivals to Elimäki across both time and space. For instance, it is possible to quantify annual departures (Mean = 135 departures ±$\pm 59$ s.d., Figure 10) and arrivals to the parish (Mean = 100 ±$\pm 41$ s.d., Figure 10). It is worth noting that there are no records of arrivals to Elimäki for the years 1914 and 1915. This is due to the year identification failing in few pages, and the records from these years are merged with records from 1916 and 1917. The mean number of departures and arrivals every year corresponds to an average of 3% and 2% of the population in Elimäki (based on a census record from 1875 giving a population of 5,305 inhabitants). These rates of migration are lower than those found in 19th century Finland using a genealogical dataset (14% and 20%, (Nitsch et al., 2023, 2016)) and in other Western historical populations (16% in historical Germany (Beise and Voland, 2008) and 36% in historical Sweden (Clarke and Low, 1992)). The number of yearly departures and arrivals is stable through time, which is unexpected considering the development of the railway (Alvarez-Palau and Martí-Henneberg, 2020) and the general increase of migration in Finland and neighbouring countries throughout the period (Svalestuen, 1977; Clarke and Low, 1992).
@@
Spatially, the initial overview of the destinations and the origins of migration shows that most movements were local (see Figure 11 and Supplementary videos available on the project repository [^3] https://github.com/TurkuNLP/finnish-migration-data for a yearly overview), in line with patterns highlighted in other studies in Finland (Nitsch et al., 2016) or other European countries from the same period (Clarke and Low, 1992; Beise and Voland, 2008).
@@
Differences in patterns of migration between populations may arise from disparities in historical sources (e.g. using military records (Kesztenbaum, 2008) or immigration registries (Engman, 1978)), in the spatial scale considered (e.g. national vs local scale), or in inclusion/exclusion criteria (e.g. only including individuals with information across their entire life, (Nitsch et al., 2023)). Therefore, interpreting differences in migration patterns between studies should be conducted with extreme caution. As this example contains information only from a single parish, the patterns described here are unlikely to be representative of the general migration patterns at the national scale.
@@ 6 Conclusions and future work
This work demonstrates that large-scale, automated extraction of structured data from handwritten historical records is feasible, even though it remains technically challenging. Using digitized Finnish moving records, we created a dataset containing over six million entries from approximately 200,000 images dated between the 19th and early 20th centuries. These entries document individual and family migration between parishes and provide a valuable source for demographic and historical analysis. The resulting dataset has a potential to support studies in many directions. First, it could be used to study migration per se, including spatial and temporal trends (e.g. changes from short to long distance migration with development of railway in Finland (Alvarez-Palau and Martí-Henneberg, 2020)), individual’s drivers of migration (age, family structure) or the impact of specific historical events on migration patterns (e.g the Finnish famine of 1866–1868, urbanization). For instance, a previous study using Russian registries documents a strong pattern of migration from eastern Finland to St-Petersburg during the 19th century (Engman, 1978), and this dataset could shed light on the scale of this migration, on the origin or the profile of migrants (age or socio-economic status). Second, migration is a key determinant of the population structure and has impacts on many characteristics of the population (e.g. population genetic structure, demographic structure). Therefore, using this dataset in combination with existing ones on Finland (records of births, deaths, past epidemics, language variation, population genetic structure) would enable accounting for and bring insights on the role of migration in explaining several processes, such as cultural transmission (e.g. of language variants (Honkola et al., 2018), disease transmission (Nitsch et al., 2025), or changes in genetic structure (Kerminen et al., 2021)). In a case study based on the Elimäki parish, we illustrated how migration volumes and patterns can be reconstructed locally.
@@
Our pipeline combines deep learning methods for preprocessing, layout analysis, table detection, cell classification, and handwriting recognition. Its modular structure allows each stage to be improved independently. While we have already demonstrated the extraction output to be sufficient for certain quantitative applications, there are still challenges that need to be addressed in future development, as shown by the error analysis and the case study. First, we aim to enhance the extraction of document metadata, such as table type, table titles, and column headers. Additional metadata is needed for conducting studies similar to our case study, especially when working with handdrawn books, where the expected columns on a page are not known in advance. To address this, we intend to investigate several options; in cases where headers are included in a page, we plan to build a specialized recognition model for extracting this information. In addition, especially for cases where headers are not included, we will investigate using LLM-based column header inference. This can be approached from two directions; a language model can be asked to inspect the extracted data table and infer a header for each column (e.g. date, name, and parish), or a language model can be requested to infer the most likely column for the given header (e.g. which column includes the parish information).
@@
Second, we are developing LLM-based postprocessing tools to standardize variations in the extracted data. This will help to normalize e.g. names, places, and dates including both orthographic variation as well as misread letters into standardized spelling, as well as infer missing values where contextual patterns in the data suggest repetition or shared information. For example, the migration records in a book should be ordered by date, and therefore the order can help us to infer missing or unclear date information.
@@ Acknowledgements
Data Accessibility
The dataset used in this study is openly available via Zenodo at https://zenodo.org/record/15606656 under the CC-BY 4.0 license. The dataset has been prepared in accordance with the FAIR principles and includes structured data in open, non-proprietary formats (e.g., XML, CSV), along with accompanying documentation to ensure reuse and transparency.
To support ongoing development and access to updated versions of the data, we also maintain a project repository at https://github.com/TurkuNLP/finnish-migration-data, which links to the corresponding Zenodo releases. The data contains no personal or sensitive information and is derived from publicly available historical records provided by the Finnish Family History Association (FFHA), whose publicly available material can be used, copied, and linked freely.